OpenStudy (anonymous):

Quadratic Formulas - 10 questions tried 6 times and failed epically. just don't get it. 1) When the sum of 5 and twice a positive number is subtracted from the square of the number, 0 results. Find the number. 2) 2/5x^2 -x+ 19/15 =0 3) .6x^2 = 1.2x +0.6 4) A rule for estimating the number of board feet of lumber that can be cut from a log depends on the diameter of the log. To find the diameter required to get 121 board ft of lumber, the equation (d-4/4)^2 = 121. are both answers reasonable? 5) 9x^2 -7x -5=0 6) 5x^2 - 2x +13 = 18x +1 7) 6p^2 = -19p -3 8) 5x^2 = -20x

OpenStudy (anonymous):

9) I can do this one :) 10) 5x^2 = 3 +2x

OpenStudy (hybrik):

Do you know the quadratic formula?

OpenStudy (anonymous):

ax2 + bx +c = 0

OpenStudy (anonymous):

x^2

OpenStudy (hybrik):

You have to put it in that form for the quadratic formula.

OpenStudy (anonymous):

I did 8&9

OpenStudy (anonymous):

oh the x = -b (sqrt b-4(a)(c))/ 2(a)

OpenStudy (hybrik):

Yes, \[x =\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a}\]

OpenStudy (hybrik):

you have to plug in the points.

OpenStudy (hybrik):

Try #1.

OpenStudy (anonymous):

x^2 + 2x+5=0 ?

OpenStudy (hybrik):

To do this, you can plug in the points to get:

OpenStudy (hybrik):

wait keep in mind if it is just x^2, the value is 1

OpenStudy (hybrik):

Okay so you would have \[x=\frac{ -2\pm \sqrt{(2)^2-4(1)(5)} }{ 2(1) }\]

OpenStudy (anonymous):

I gtg, will be on tomorrow can you help then?

OpenStudy (anonymous):

no available computer till tomorrow

OpenStudy (hybrik):

Okay.

OpenStudy (anonymous):

Thank You so much. Very appreciated

OpenStudy (hybrik):

Your welcome.

OpenStudy (anonymous):

um...so did you still want to help me? lol

pooja195 (pooja195):

Can you make a post for each question so its easier to help? .-.

OpenStudy (anonymous):

I mean, yeah I can lol.

pooja195 (pooja195):

Thanks :3